Please help!

Solve by elimination method:

4x+ 5y = 2
8x + 10y = 4
Either the solution is an ordered pair (____,____)
or
there are infinitely many solutions ____

Answer by checkley77(12844)

4x+5y=2 MULTIPLY THIS EQUATION BY -2 & ADD
8x+10y=4
-8X-10Y=-4
-------------------
0X+0Y=0
There are infinitely many solutions (8,-6), (0,0.4), (-3,2)